OpenStudy (anonymous):

the sum of two numbers is 110. Their difference is 28. What is the smallest of the two numbers? a.56 b.41 c.69 d.55

OpenStudy (anonymous):

x+y=110 x-y=28 Can you see it now?

OpenStudy (anonymous):

So you have two variable a and b that follow the following equations: a+b = 110 and a-b = 28 Therefore, you can say (a + b) - (a - b) = 110 - 28 Work this out and you'll get the answer.

OpenStudy (anonymous):

i got 82

OpenStudy (anonymous):

110-41=69 69-41=28

OpenStudy (anonymous):

41 is the correct one

OpenStudy (anonymous):

b

Parth (parthkohli):

x + y = 110 x - y = 28 2x = 138 x = 69 So, 69 and 69-28 = 41 are answers

Parth (parthkohli):

41 is the final answer

OpenStudy (anonymous):

thanks everyone!

Parth (parthkohli):

^^You're welcome :D

OpenStudy (anonymous):

\[\left\{ x+y=110; x-y=28 \right\}\] Very simple.